# LLM Analysis of 150+ years of German Parliamentary Debates on Migration Reveals Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

### v1

@@ 4 Temporal Trends in Solidarity Towards Migrants
First, we conduct a stability test by randomly sampling 200 subsets of at least 5 keywords each and comparing the resulting high-level trends using Pearson’s correlation coefficient. Across all 19,900 pairwise comparisons, the trends are highly consistent (average correlation of 0.92 for solidarity, 0.81 for anti-solidarity, and 0.94 for mixed; with 25% quantiles of 0.90, 0.70, and 0.92, respectively). [^6] For more detail on keywords and the stability test, see A. Second, we restrict the analysis to the two most frequent and temporally widespread keywords, Ausländer (en.: foreigners) and Flüchtlinge (en.: refugees), which occur in every decade (19.7k and 12k mentions, respectively). Together, the resulting plots (Figure 23 in the Appendix) and the stability test results confirm that overall trends in solidarity and anti-solidarity are preserved both under this restricted vocabulary and across different keyword selections.
4.1.2 Notable Trends in 1867-2022.
Inspecting the trends in Figure 7, we observe four periods with notable shifts in solidarity: Between 1890-1920, in the 1930s, in the post-war period, and in the recent past after 2010.
Regarding the 1890-1920 period, we observe some disagreement between the models: GPT-4 shows an increase in solidarity between 1900 and 1920, while Llama-3.3-70B does not. We also note that the agreement between human annotators and between models and annotators is lower for historic data, such that we do not consider this trend to be sufficiently reliable for further analysis.
The 1930s and 1940s data points are outliers, as the parliament was swiftly sidelined by the Nazi regime and the only 109 available samples from 1933 to 1949 mostly concern inter-state relations (mentioning, e.g., ?foreign creditors?) that fall into the ?None? category. Hence, we exclude these data points from further analysis, as well.
For the two remaining periods, 1949–1957 and 2009–2025, we take a more detailed look. In the postwar years (Section 4.2), parliamentary discourse reveals high levels of solidarity and little anti-solidarity toward migrants, despite the severe economic and infrastructural strain of the time braun2014immigration (14). This raises our first guiding question: Why is there so much solidarity (and so little anti-solidarity), and what was it based on? In contrast, the recent period (Section 4.3) shows a clear rise in anti-solidarity and a decline in solidarity, beginning in the early 2010s. These shifts likely reflect the political impact of recent refugee movements and the growing influence of anti-immigration views, such as those promoted by the far-right party Alternative für Deutschland (AfD) (en.: Alternative for Germany), which has gained substantial electoral support in Germany since its founding in 2013. This leads to our second question: What are the drivers of anti-solidarity in the contemporary period?
@@ 4.2 Germany after the Second World War, 1949-1957
The Second World War set off the largest wave of forced migrations in European history, with tens of millions of people, both non-Germans and ethnic Germans, displaced or resettled across Central and Eastern Europe schulze2017commemoration (90). The populations displaced by the war were highly heterogeneous, including refugees, displaced persons (DPs), forced laborers, and—forming the largest group of expellees after 1945 in both Europe as a whole and Germany—ethnic Germans. For those who settled in postwar Germany, integration posed significant political and societal challenges, shaped by regional disparities, economic hardship, and tensions with the local population chevalier2024forced (21).
This subsection focuses on the early postwar period (1949–1957), when the Federal Republic of Germany institutionalized solidarity as part of nation-building. This period is analytically important because solidarity was no longer merely a moral gesture, but a tool of governance, embedded in legal frameworks like the Lastenausgleich (LAG) (en.: Equalization of Burdens), which redistributed resources between the war-damaged and the less affected to promote social cohesion and democratic legitimacy kittel2020stiefkinder (54). Crucially, such measures applied specifically to Volkszugehörige (en.: ethnic Germans), excluding other migrant groups from the benefits provided by these laws. However, since solidarity in this period was neither static nor uniform, shaped by diverse migrant groups, competing political ideologies, and contested notions of deservingness ther1996integration (96), we draw on Thijssen’s typology of solidarity thijssen2012mechanical (97) to guide our analysis. By distinguishing between identity-, need-, and contribution-based forms, this framework allows us to examine how solidarity was constructed and sustained amid institutional strain.
@@ 4.2.1 Overall solidarity distribution
Throughout the 1949–1957 period, solidarity dominated parliamentary discourse on postwar population movements. As shown in Figure 9, [^7] We use raw counts as they provide a more interpretable view of stance distribution during this period, given the overwhelming dominance of solidarity and the low frequency of categories like anti-solidarity. mentions of solidarity consistently outpaced both anti-solidarity and mixed statements, at times reaching ten times their frequency.
@@
Figure 9: Distribution of (Anti-)solidarity mentions in Bundestag debates, 1949–1957 (raw counts). LAG refers to the Lastenausgleichsgesetz (en.: Equalisation of Burdens Law); BVFG to the Bundesvertriebenengesetz (en.: Federal Expellees Act); HAuslG to the Gesetz über die Rechtsstellung heimatloser Ausländer im Bundesgebiet (en.: Law on the Legal Status of Stateless Foreigners in the Federal Territory).
@@
Mentions of solidarity closely track key legislative moments, particularly in the debates leading up to the passage of major laws. A major spike in March 1950 aligns with early discussions on the integration of expellees and Article 131, which addressed the rehabilitation of former public servants of the Nazi regime, many of them displaced. A second cluster from January to May 1952 reflects debates on the LAG, preceding its adoption in May (see an excerpt from such a debate in Example 1 in Table 10 in the Appendix). Another sharp rise in February–March 1953 corresponds to readings of the Bundesvertriebenengesetz (BVFG) (en.: Federal Expellee Law), which defined the legal status of expellees and refugees and codified rights and integration measures, see Example 2 for such a debate. Later peaks in September 1954 and June 1955 reflect continued debate over LAG implementation and budget issues. Smaller but sustained increases in 1956 relate to new support measures for Sowjetzonenflüchtlinge (en.: Soviet-zone refugees) and Aussiedler (en.: resettlers).
@@
Figure 10: Distribution of (Anti-)solidarity frame subtypes from 1949 to 1957 (raw counts). An inset plot shows the evolving composition of solidarity subtypes, normalized within each 6-month period. Values represent the proportional share of each subtype among all solidarity mentions during that period.
The prominence of solidarity in this period is unsurprising. Parliamentary debates used it not only to justify compensation, but also to navigate postwar fragmentation and competing claims to victimhood and belonging hughes1999shouldering (49). This multifaceted nature of solidarity is evident in the distribution of all four subtypes across key legislative moments, as shown in Figure 10. Group-based and compassionate frames were especially prominent, reflecting an emphasis on both shared national belonging and moral obligation toward displaced populations (see, e.g., Example 1 in Table 10, where solidarity is tied to the right to the homeland and framed as a legal entitlement). More striking is the shift in the balance between these two subtypes over time. As shown in the inset of Figure 10, group-based solidarity dominated in the early 1950s but declined after mid-1953, while compassionate steadily increased. This suggests a rhetorical transition: once legal belonging was formalized, most notably with the LAG in 1952 and BVFG in 1953, solidarity no longer needed to be framed primarily in terms of shared identity. The shift may also reflect demographic change within the displaced populations, as newer arrivals like Sowjetzonenflüchtlinge from the GDR and Aussiedler (en.: resettlers) from the Eastern Bloc countries were less easily framed as co-nationals boke1996fluchtlinge (13). This trend aligns with Thijssen’s reading of Durkheim: as group-based solidarity extends to more diverse groups, it gives way to need-based appeals thijssen2012mechanical (97) (see Example 3 in Table 10 in the Appendix).
These patterns show that solidarity was used in parliamentary debates to manage the challenges of postwar reconstruction, as reflected in key legislative reforms. While the claimant groups of the broadest measure LAG were highly diverse (bombed out, evacuated, expelled, financially damaged), their eligibility for compensation under the law was based on Volkszugehörigkeit (en.: German ethnicity), excluding many from formal recognition kittel2020stiefkinder (54, 73). The concept of Volksdeutsche was a fluid and constructed Nazi category whose ambiguous definition nonetheless had concrete and deadly implications bergen2005tenuousness (9). We therefore examine how (anti-)solidarity subtypes were expressed across key migrant-related terms in parliamentary discourse in that period.
@@ 4.2.2 Solidarity Frames Across Keywords in 1949-1957
This analysis includes only migrant-related keywords that appear at least 200 times in combination with a solidarity, anti-solidarity subtype or mixed stance. Keywords dominated by none or with low total counts are excluded (e.g., Asylsuchende (en.: asylum seekers), Einwanderer (en.: immigrants), Emigranten (en.: emigrants), etc.); see Table 5 for full distributions.
Figure 11 shows the distribution of (anti-)solidarity frames across migrant-related keywords in 1949–1957. Across all groups, solidarity was far more common than anti-solidarity, with group-based and compassionate frames dominating. However, the distribution of subtypes varies between groups.
@@
Vertriebene (en.: expellees) are most often associated with group-based solidarity, accounting for around half of all expressions with a stance, while also receiving a substantial share of compassionate solidarity (nearly 30%). This reflects their dual position: as key representatives of German victimhood (alongside prisoners of war) during and after the World War II moeller2001war (71), they were both integrated into West Germany’s nation-building narrative and portrayed as victims of Allied or foreign injustice. Still, around 5–6% of stance expressions toward Heimatvertriebene fall into the anti-solidarity category, particularly group-based (solidarity limited to ?us?, the in-group). These expressed concerns that expellees strained local infrastructures and were unfairly privileged over equally suffering Einheimische (en.: locals), mirroring broader public resentment kossert2008kalte (56). Perceptions of injustice were heightened by the uneven distribution of expellees across West German states braun2014employment (15), particularly in regions where their prominent presence in public institutions and politics sparked fears of Überfremdung (en.: overforeignization) and contributed to a sense of unequal burden-sharing levy2005memories (64) see Example 4 in Table 10.
@@
Figure 11: Relative distribution of (Anti-)Solidarity subtypes across selected migrant-related keywords, 1949–1957. Bars represent the proportion of each subtype within each keyword group, thus all subtypes for a given group sum to 100%.
@@
Flüchtlinge (en.: refugees), Sowjetzonenflüchtlinge (en.: Soviet-zone refugees), and Aussiedler (en.: resettlers) are primarily framed through compassionate solidarity (45–52%), followed by group-based solidarity (25–30%). The label Flüchtlinge initially functioned as a broad catch-all category in the late 1940s nachum2018semantics (74). With the 1953 BVFG, it was more clearly differentiated into Vertriebene, Sowjetzonenflüchtlinge, and Aussiedler. Yet their Germanness was not valued equally: while Vertriebene were framed as victims of expulsion and central to West German nation-building, the label Flüchtlinge retained post-war connotations of marginality and foreignness nachum2018semantics (74). This helps explain a higher share of compassionate solidarity, as well as the notable anti-solidarity (around 10%). Sowjetzonenflüchtlinge, in turn, were primarily framed as Cold War refugees fleeing communist oppression and thus not integrated into the nation building narrative like the Vertriebene, prompting more compassionate than group-based solidarity nachum2018semantics (74, 13) (see Example 5).
@@
Ausländer (en.: foreigners) were the main targets of anti-solidarity. Not a single coherent group, they were a discursively constructed ?other?, including DPs, Eastern European border crossers, wartime laborers, and early foreign workers smith1994germany (93). Often contrasted with German groups like Vertriebene or Einheimische, they were framed as non-belonging, economically burdensome, and sometimes criminalized holian1945missing (44) (see Example 6). However, despite this exclusionary framing, solidarity with Ausländer, especially group-based, exceeded anti-solidarity. This might be partly due to the category’s fluidity, which included ethnic Germans not yet formally recognized as such ruhkopf2023institutionalisierte (86) in the early postwar years. Politicians like Lukaschek also linked DPs and expellees through shared narratives of lost Heimat (en.: homeland), enabling symbolic inclusion into the national victim community ruhkopf2023institutionalisierte (86) (Example 7). The second-highest solidarity subtype, exchange-based solidarity, reflects the perceived political utility of Ausländer in strengthening West Germany’s international legitimacy within the Western alliance ruhkopf2023institutionalisierte (86) (Example 8). The 1951 Homeless Foreigners Law (HAuslG), passed under Allied pressure, granted DPs civic and social rights largely to meet conditions for sovereignty and improve the country’s international standing holian1945missing (44, 86). At the same time, instances of genuine solidarity toward non-ethnic German foreigners, particularly in CDU and SPD, statements, show that solidarity with Ausländer was not limited to Germanness or foreign policy utility but also reflected broader humanitarian commitments (see Examples 7 and 9).
@@
In sum, solidarity was expressed differently depending on the migrant group and was shaped by perceived ?deservingness?. Vertriebene were largely included as part of the national community, while others, especially Ausländer, were treated as outsiders and received comparatively less support. These differences reflected not only group characteristics but also the ideological strategies of political parties. The next subsection examines how different parties in the early Bundestag framed (anti-)solidarity in their parliamentary speech.

### v2

@@ 4.1.1 Robustness to Keyword Composition
First, we conduct a stability test by randomly sampling 200 subsets of at least five keyword groups (grouping inflectional variants). For each subset, we re-estimate the high-level DSL trends (as defined in Section˜3.2.2) and compare the resulting time series to the full-sample trend using Pearson’s correlation coefficient. The trends remain broadly stable, with mean correlations of 0.93 for solidarity, 0.88 for anti-solidarity, and 0.82 for mixed; for more detail on keywords and the stability test, see Appendix˜A in the Appendix. Second, we restrict the corpus to the two most frequent and temporally widespread terms, Ausländer (en.: foreigners) (19.7k mentions) and Flüchtlinge (en.: refugees) (12k), and re-estimate DSL on this reduced sample. Both checks confirm that overall solidarity and anti-solidarity trends are preserved both under this restricted vocabulary (see Figure˜33) and across different keyword selections.
4.1.2 Notable Trends in 1867-2025
Inspecting Figure˜10, we identify three periods with notable shifts in solidarity: 1890-1900s, the post-war period, and the recent period after 2010. We note that the 1930s and 1940s are outliers, as the parliament was swiftly sidelined by the Nazi regime and the corpus contains only 109 available instances from 1933 to 1949. Hence, we exclude these data points from further analysis.
The first of these, 1890-1900, is difficult to interpret with confidence. In this early part of the series, even the strongest model in our ensemble performs worse than in later decades in terms of macro-F1, and agreement between human annotators is also lower (see Figure˜18). Although DSL corrects the resulting estimates, we do not consider this early trend sufficiently reliable for detailed substantive analysis. This period also contains fewer instances than the post-war and recent periods that we analyze in detail below (see Figure˜2).
For the two remaining periods, 1949–1957 and 2009–2025, we take a more detailed look.
@@ 4.2 Germany after the Second World War, 1949-1957
The Second World War triggered the largest wave of forced migration in European history, with tens of millions of people displaced or resettled across Central and Eastern Europe (Wessel, 2017). The affected populations were heterogeneous—refugees, displaced persons (DPs), forced laborers—and—forming the largest group of expellees after 1945 in both Europe as a whole and Germany—ethnic Germans. For those who settled in postwar Germany, integration posed political and societal challenges shaped by regional disparities, economic hardship, and tensions with locals (Chevalier et al., 2024).
This subsection examines the early postwar period (1949-1957), when the Federal Republic institutionalized solidarity as part of nation-building. This period is analytically important because solidarity became a tool of governance, embedded in laws like the Lastenausgleich (LAG) (en.: Equalization of Burdens), which redistributed resources between the war-damaged and the less affected to promote social cohesion and democratic legitimacy (Kittel, 2020). These measures applied specifically to Volkszugehörige (en.: ethnic Germans), excluding other migrant groups. Since solidarity was neither static nor uniform, shaped by diverse migrant groups, competing ideologies, and contested notions of deservingness (Ther, 1996), we employ Thijssen’s typology of solidarity (Thijssen, 2012) to guide our analysis. By distinguishing identity-, need-, and contribution-based forms, we analyze how solidarity was constructed and sustained amid institutional strain.
@@ 4.2.1 Overall solidarity distribution
During 1949–1957, solidarity dominated parliamentary discourse on postwar population movements. As Figure˜11 shows, [^7] We report (DSL-adjusted) counts rather than proportions (see definition (d) in Section 4.2.1) because counts make spikes visible whenever migration-related laws are debated and enable comparison of debate volume across months. mentions of solidarity consistently exceeded anti-solidarity and mixed statements, at times by a factor of ten.
@@
Figure 11: Estimated counts of (anti-)solidarity mentions in Bundestag debates, 1949–1957 (DSL-adjusted, see definition (d) in Section˜4.2.1). LAG == Lastenausgleichsgesetz (en.: Equalisation of Burdens Law); BVFG == Bundesvertriebenengesetz (en.: Federal Expellees Act); HAuslG == Gesetz über die Rechtsstellung heimatloser Ausländer im Bundesgebiet (en.: Law on the Legal Status of Stateless Foreigners in the Federal Territory).
@@
Mentions of solidarity closely track key legislative moments. A major spike in March 1950 aligns with early discussions on the integration of expellees and Article 131, which addressed the rehabilitation of former public servants of the Nazi regime, many of them displaced. A second cluster in early 1952 reflects debates on the LAG, preceding its adoption in May (see Example 15 in Table˜19 in the Appendix). Another sharp rise in February–March 1953 corresponds to readings of the Bundesvertriebenengesetz (BVFG) (en.: Federal Expellee Law), which defined the legal status of expellees and refugees and codified integration measures (Example 15 ). Later peaks in September 1954 and June 1955 reflect debate over LAG implementation and budgets. Smaller but sustained increases in 1956 relate to new support measures for Sowjetzonenflüchtlinge (en.: Soviet-zone refugees) and Aussiedler (en.: resettlers).
Box 4.1: Estimated Monthly Counts of (Anti-)Solidarity Speeches are grouped by month $t$, with $N_{t}$ speeches in month $t$. We define the estimated number of mentions of label $\ell$ in month $t$ as: (d) DSL-adjusted monthly counts: $\hat{C}^{DSL}_{t,\ell}=\sum_{i=1}^{N_{t}}\tilde{Y}_{i,\ell}$, where $\tilde{Y}_{i,\ell}$ denotes the DSL-adjusted class score for label $\ell$ in speech $i$, as defined in Equation˜1 in Section˜2.
@@
Figure 12: Distribution of (anti-)solidarity subtypes, 1949–1957 (DSL-adjusted expected counts, see definition (d) in Section˜4.2.1). Inset shows the evolving composition of solidarity subtypes, normalized within each six-month period as shares of all solidarity framing (see definition (c) in Section˜3.2.2).
The prominence of solidarity in this period can be explained by the fact that debates used it not only to justify compensation but also to manage postwar social and economic fragmentation and competing claims to belonging (Hughes, 1999). The distribution of all four subtypes across key legislative moments (Figure˜12) reflects this mix of solidarity rationales. Compassionate and group-based frames dominate throughout the period (see inset of Figure˜12). By emphasizing shared national belonging (group-based) and the suffering of affected populations (compassionate), these frames construct a victim-centered Schicksalsgemeinschaft (en.: community of fate), uniting Germans through shared loss (Gengler, 2020) (see Example 15in Table˜19 in the Appendix [^8] Example labels are based on the dominant DSL subtype, that is, the subtype with the highest DSL-adjusted score. This is useful for illustrating which subtype is most strongly reflected in a given speech within the DSL-based analysis; however, since DSL is designed for aggregate inference, these labels do not guarantee corrected true labels for individual instances.). This coexistence of the two subtypes is consistent with Thijssen (2012), who argues that compassionate and group-based solidarity form a dialectical pair: identification with a collective (?we?) generates obligations toward its members, while emphasizing their suffering reinforces that collective bond.
Additionally, a small but persistent share of exchange-based solidarity appears throughout the period. It reflects debates over efficient allocation and the reintegration of displaced populations into labor and housing markets, with measures such as the LAG framed as tools to rebuild productive capacity (Examples 15 and 18).
These patterns show that solidarity in parliamentary debates mediated the challenges of postwar reconstruction through key legislative reforms. However, since these reforms differentiated between migrant groups, we examine how (anti-)solidarity subtypes were expressed across key migrant-related terms in this period.
@@ 4.2.2 Solidarity Frames Across Keywords in 1949-1957
The analysis includes migrant-related keywords for which the DSL-adjusted expected number of stanced instances (i.e., solidarity, anti-solidarity, or mixed rather than none) is at least 150. Keywords dominated by none or with low DSL-adjusted expected counts (e.g., Asylsuchende (en.: asylum seekers), Einwanderer (en.: immigrants), Emigranten (en.: emigrants)) are excluded; see Table˜9 in the Appendix for full distributions. Figure˜13 shows the distribution of (anti-)solidarity frames across these terms.
Vertriebene (en.: expellees), a category reserved primarily for Volkszugehörige (en.: ethnic Germans), are most often associated with compassionate solidarity (about half of all expressions with a stance), and also with group-based solidarity (nearly 25%). This reflects their dual position: as key representatives of German victimhood during and after the World War II (Moeller, 2001), they were both central to West Germany’s nation-building narrative and portrayed as victims of Allied or foreign injustice.
Box 4.2: Distribution of (Anti-)Solidarity Subtypes Across Groups Speeches are grouped by a categorical variable $g$ (a keyword or a political party). $I_{g}$ is the set of speeches belonging to group $g$. The share of subtype $\ell$ within group $g$ is defined as: (e) Subtype shares within groups: $\hat{s}_{g,\ell}=\frac{\hat{M}_{g,\ell}}{\hat{M}_{g}}$, where $\hat{M}_{g,\ell}=\sum_{i\in I_{g}}\tilde{Y}_{i,\ell}$ is the DSL-adjusted mass of subtype $\ell$ in group $g$, where $\tilde{Y}_{i,\ell}$ denotes the DSL-adjusted class score for label $\ell$ in speech $i$, as defined in Equation˜1 in Section˜2. $\hat{M}_{g}=\sum_{\ell\in L}\hat{M}_{g,\ell}$ is the total mass of group $g$ across all labels $L$ (four solidarity subtypes, four anti-solidarity subtypes, and mixed stance). Shares sum to one within each group $g$.
@@
Still, around 10% of stance expressions toward Heimatvertriebene were anti-solidarity, particularly group-based (solidarity limited to ?us?, the in-group), reflecting concerns that expellees strained local infrastructures or were unfairly privileged over equally suffering Einheimische (en.: locals), mirroring broader public resentment (Kossert, 2008). Perceptions of injustice grew with the uneven distribution of expellees across West German states (Braun and Mahmoud, 2014), especially where their prominence in institutions and politics sparked fears of Überfremdung (en.: overforeignization) and unequal burden-sharing (Levy and Sznaider, 2005) (see Example 15 in Table˜19).
@@
Figure 13: Relative distribution of (anti-)solidarity subtypes across selected migrant-related keywords, 1949–1957 (DSL-adjusted estimates; see definition (e) in Section˜4.2.2). Bars show each subtype’s share within a keyword group; shares sum to 100%. Full label distributions appear in Table˜9 in the Appendix.
@@
Flüchtlinge (en.: refugees), Sowjetzonenflüchtlinge (en.: Soviet-zone refugees), and Aussiedler (en.: resettlers) are also primarily framed through compassionate solidarity (50–72%), followed by group-based solidarity (15%). The label Flüchtlinge began as a broad catch-all in the late 1940s but, even after the 1953 BVFG differentiated it into Vertriebene, Sowjetzonenflüchtlinge, and Aussiedler, it retained connotations of marginality and foreignness (Nachum and Schaefer, 2018). This helps explain a more notable share of anti-solidarity (around 10%). Sowjetzonenflüchtlinge, framed as Cold War refugees from communist oppression, were less integrated into the nation-building narrative than Vertriebene, prompting more compassionate than group-based solidarity (Nachum and Schaefer, 2018; Böke, 1996) (see Example 16).
@@
Among the migrant categories, Ausländer (en.: foreigners) shows one of the highest relative shares of anti-solidarity. A discursively constructed ?other?, including DPs [^9] With the 1951 HAuslG, authority over the remaining DPs was transferred from Allied administration to the Federal Republic, and they were reclassified as heimatlose Ausländer (en.: stateless foreigners), a noncitizen status that granted them most civil rights without granting citizenship (Holian, 2011)., Eastern European border crossers, wartime laborers, and early foreign workers (Smith, 1994), they were contrasted with groups like Vertriebene or Einheimische and framed as non-belonging (evident in group-based anti-solidarity share), economically burdensome (exchange-based anti-solidarity), or even criminalized (Holian and Wilhelm, 2017) (see Example 16). Yet solidarity, especially group-based, still exceeded anti-solidarity, partly due to the category’s fluidity: it sometimes included ethnic Germans not yet formally recognized as such in the early postwar years (Ruhkopf, 2023). Speakers reinforced this dynamic by linking DPs and expellees through shared narratives of lost Heimat (en.: homeland), enabling symbolic inclusion into the national victim community (Ruhkopf, 2023) (Example 16).
The second-highest subtype, exchange-based solidarity, reflects the perceived political utility of Ausländer in strengthening West Germany’s international legitimacy (Ruhkopf, 2023), for example through the 1951 Homeless Foreigners Law (HAuslG), passed under Allied pressure (Holian and Wilhelm, 2017) (Example 17). However, solidarity also extended to non-ethnic German foreigners, particularly in CDU and SPD statements, showing that solidarity with Ausländer was not limited to Germanness or foreign policy utility but also reflected broader humanitarian commitments (see Examples 16 and 17).
@@
In sum, compassionate framing dominated across most categories, reflecting a broader discourse of shared suffering in which displaced populations were positioned as victims of the war and its consequences. Nevertheless, solidarity was still shaped by perceived ?deservingness? and ethnic identity. While Vertriebene were more firmly incorporated into the national community as ethnic Germans, others, especially Ausländer, received comparatively less support and conditional inclusion (Nachum, 2013). These differences reflected both group characteristics and the ideological strategies of political parties. The next subsection examines how different parties in the early Bundestag framed (anti-)solidarity in their parliamentary speech.